Evaluate the definite integral by the limit definition. + 2) dx
Answer #1

Page ① 6- | (225+2) du a=-2 b=1 a-2 1-1-2) Ax= b-a h 11 h Xi = ation=-2 +1 ( 2 ) = -2+ 31 +2 h2+2 h2+2 fx) = 2x²+2 fini)=2x1²Page 2 30 - 72 ½ i + 54 i=1 h2 i=1 hyo 52 ) - 3 (930) + 5% (n=2x2 +1 = hin 30- 36(1+) + 9 (2+ +) (2+)] U dim 400 hyoo 30 - 36